OpenStudy (anonymous):

Solve -63x + 7y + 49 = 0 for y

OpenStudy (radar):

There are several ways to go about this. All will attempt to isolate the y so that it ends up by itself, and positive on the left of the equal sign.

OpenStudy (anonymous):

\[-63x+7y+49=0\] \[7y=63x-49\] \[(7y=63x-49)\div7\] \[y=9x-7\]

OpenStudy (anonymous):

@kas9987 its right! thank you

OpenStudy (radar):

Using algebra, I would first divide all terms by 7. This would result in: -9x + y + 7 = 0 Next I would add 9x -7 to both sides of the equal sign, getting y=9x-7. There you have it, y by itself on the left and 9x -7 on the right.

OpenStudy (anonymous):

@mariruiz no problem